A small glider is placed against a compressed spring at the bottom of an air track that slopes upward at an angle of 46.0 ∘ above the horizontal. The glider has mass 9.00×10−2 kg . The spring has 670 N/m and negligible mass. When the spring is released, the glider travels a maximum distance of 1.80 m along the air track before sliding back down. Before reaching this maximum distance, the glider loses contact with the spring. the distance of the spring was originally compressed is x = 5.84×10−2 What is the kinetic energy of the glider at this 0.8m?
Answer #1

Here W_{total} = W _{spring} + W _{gravity}

W_{total}  = 1/2 m v^{2} - 0 = final kinetic energy - initial kinetic energy = final kinetic energy, as initial kinetic energy = 0

here we have to understand that Work done by spring is counteracted by the work done by gravity.

\therefore   W_{total} = W _{spring} - W_{ gravity}

W_{ gravity} = mg L cos\theta where m = mass of the glider = 9.00×10−2 kg

g = Acceleration due to gravity = 9.8 m / s^{2}

L = maximum distance travelled by the glider in contact with the spring= 0.8m

=  9.00×10−2 \times 9.8 \times 0.8 \times  cos ( 90-46)

= 70.56 \times 10−2 \times cos 54

= 7056 \times10^{-4}\times 0.58

= 4092.48 \times10^{-4} J

  W _{spring} = 1/2 K x ^{2}   where K = spring constant = 670 N/m

and the distance of the spring originally compressed is x = 5.84×10−2 m

= 1 / 2 \times 670 \times 5.84× 5.84×10^{-4} = 335 \times 5.84× 5.84×10^{-4}

= 11425.76 ×10^{-4} J

\therefore   W_{total} = W _{spring} - W_{ gravity}

= 11425.76 ×10^{-4} -   4092.48 \times10^{-4} = 7333.28 \times10^{-4} J

\therefore  the kinetic energy of the glider = 7333.28 \times10^{-4} J